FUSE Mission Status Report

Mission Status Report #12      Star Date: July 30, 1999

Mission Status Reports on the web page will occur when news-worthy events occur (but at least weekly!) for the time being.

-- Bill Blair

Image: Portion of "first light" image with the FES-A Guide Star Camera.

(Click image to see full field of view version.)

(See links at bottom of page for more information.)

FUSE In-Orbit Checkout Continues

The FUSE satellite remains healthy and the In-Orbit Checkout phase is still in progress. Thermal control still has the satellite in "bake-out" mode to drive any trapped gases out of the structure prior to the beginning of normal operations.

We are monitoring the gas pressure in the spectrograph by performing daily "sniff tests", where vacuum pumps are run that provide a current reading that can be converted to a "pressure." The current pressure is very low by earthly standards, but still not low enough that the detectors can have high voltage applied for science operations. Hence, we must simply wait for the pressure to continue its decrease until a safe level is reached. The current extrapolation indicates it may be mid-August before the pressure will drop to the necessary level.

Much of the last week has been spent performing various slewing and acquisition tests. (By that, I mean we move the telescope pointing from one position in the sky to another and try to lock on to pre-planned guide stars.) All of the pointings remain in the southern "Continuous Viewing Zone" for now as we continue to avoid pointing at the sunlit earth--see last week's report. We have had some difficulties getting the Attitude Control System (ACS) on the spacecraft to accept Fine Pointing Data (FPDs) from the instrument computer. This is necessary to achieve a condition called "guiding on known stars" that will ultimately allow us to observe science targets. Our tests have shown WHY this is occurring, and we are testing a temporary (if somewhat cumbersome) workaround, but a longer term fix to this operational difficulty is being studied. It is good to be working these kinks out of the system now, so that when we can turn on the detectors, checkout can proceed smoothly from that point. Hence, while some aspects of checkout have been delayed, we are actually ahead of the game in other areas.

All of this is still being done with only one telescope door open. Over the next week we will perform more acquisition tests (there are several different modes of acquiring targets and they all need to be checked out), and continue monitoring the pressure. After that, we will open the remaining telescope baffle doors in preparation for the focus and alignment activities.

Reported by: Bill Blair, Chief of Mission Planning

Related Links:

CSA Press Release on FES First Light Image in English and in French.

Return to the List of Status Reports